# Ensembled Direct Multi Step forecasting methodology with comparison on macroeconomic and financial data

### v1

@@ 4 Comparisons on macroeconomic and financial data
We evaluate the EDMS methodology against the Ensembled Iterative Multi-Step approach to quantify the performance gains achieved through direct forecasting. In Subsection 3.1, we describe the data used, along with the process of data gathering and preparation. Subsection 3.2 explains the methodology employed in the comparative analysis, while the final Subsection 3.3 presents and discusses the results.
@@ 4.1 Data
Our comparative analysis utilizes carefully selected, common, and publicly available datasets. We chose public data for transparency and to ensure that our methodology is tested with data that is relatively similar to that applied for practical purposes. Moreover, to ensure the objectivity of our results, we employ a well-defined process of data selection and preparation, which can be broken down into the following stages:
@@
    A set of common macroeconomic indicators is selected.
@@
    For the selected indicators, we identify the public source with the highest quality of data, specifically focusing on methodologically consistent datasets across multiple countries for the same indicators. Furthermore, countries are selected based on two primary criteria: geographical diversity and the level of economic development, aiming to capture a representative cross-section of the global economy.
@@
    For the forecasting implementation, it is essential that all time series have equal lengths. However, the time series in the dataset vary in length. We refrain from reducing all time series to the shortest one, as this would significantly decrease the size of the dataset and ultimately diminish the quality of model training. Instead, we remove only those time series with the smallest number of observations while simultaneously reducing the overall length of all time series in a manner that maximizes the number of data points in the entire dataset.
@@
    We eliminate time series that are highly irregular and cause model fit errors that are disproportionately higher than those of other time series in the specific dataset.
@@
For data description with list of specific countries see Appendix A. Additionally, further data was selected from the Connect Platform, which is available through S&P Global Market Intelligence, a trusted data provider:
@@
For data description with list of specific countries see Appendix A. In the CPI and Short Interest Rate datasets, we excluded Brazil due to a sharp increase in its values for some periods, which caused the forecast error to peak at approximately 50 times higher than the errors of all the remaining countries. A similar situation occurred for Brazil and Peru in the Wholesale Producer Price Index, where the forecast error peak was around 10 times greater than that of other countries.
Additionally, a comprehensive set of Harmonized Index of Consumer Prices (HICP) indicators was sourced from Eurostat for EU countries. These indicators included numerous concepts, all rebased to 2015. For the list of concepts see Appendix A. These indices were analysed over the period from December 2000 to December 2024, in accordance with the broader dataset constraints.
@@
  • FTSE 100 refers to a broader index of major companies listed on the London Stock Exchange, reflecting the performance of the UK market
@@
  • Nikkei 225 monitors 225 large, publicly owned companies on the Tokyo Stock Exchange and serves as a primary indicator of Japan’s stock market performance.
@@
We used daily data from date 2021.07.14 until 2025.07.14.
We divided each dataset into a training set and a testing set. The training set is used to train the models, while the testing set is employed to measure the performance of the methodology. One of the main features of DMS is its ability to avoid the accumulation of errors in consecutive forecasts, a problem that often occurs in IMS. Therefore, the testing horizon must be sufficiently long. The another reason for large testing set is that difference from IMS might be rather small in small forecasting horizons, as the first retraining is carried out after one year/month, and for that time differences in forecasts might still relatively small. After the second retraining, that is, in five years/months time, the difference should be more visible. At the same time there should be sufficient training data to train models, and the balance was found to have around 20 years/months of testing data.
@@ 4.2 Comparison methodology
We measure the performance of the EDMS methodology by comparing it with EIMS, which is a conventional forecasting methodology and serves as a natural benchmark. The comparison involves calculating the average Mean Absolute Percentage Error (MAPE) for each time series in the datasets, denoted as $\overline{mape}$. We selected MAPE because this metric is well-suited for comparing multiple time series and calculating an average metric. For instance, in the case of Mean Absolute Error (MAE), each time series may have values of different magnitudes, leading to considerable variation in MAE scores. Consequently, the contributions of some series to the overall average may distort the final value. In contrast, MAPE mitigates this issue since all errors are expressed on the same scale, that is, percentages.
@@
where $M$ represents the number of time series for a particular dataset, $H$ is the number of observations in the forecasting horizon, $y_{t+h,j}$ are the testing observations of the $j$-th time series, and $t+h$ is the time step at which forecasts are computed. Here $\widehat{y}_{t+h}^{j}$ can be either, EDMS or EIMS forecast.
Recall that the forecast in EIMS methodology, denoted as $\check{y}_{t+h}$ at time step $t+h$ can be expressed by the following formula:
$\check{y}_{t+h}=\sum_{i}w_{t,i}\widetilde{y}_{t+h,i}(Y_{t}),$
where $w_{t,i}$ are the weights computed from the initial dataset, $h=1,2,\ldots,H$ are the steps in the horizon, and $\widetilde{y}_{i,t+h}$ are the forecasts at a time step one year ahead from time $t$ for the $i$-th model, which is trained on the data up to time $t$, denoted as $Y_{t}=(y_{1},y_{2},\ldots,y_{N})$.
@@ 4.3 Results
In Tables 1, 2 and 3 we present the results of our computations for economic data, where each row corresponds to a distinct dataset. We categorize the results into annual, quarterly, and monthly frequencies due to the substantial differences in the amount of training data used for each frequency, which significantly impacts the analysis results. In addition to the average MAPE errors (4) for the EIMS and EDMS methodologies and the values of $\Delta(\%)$, we also list the number of time series, the total number of observations, and the number of observations for testing. At the bottom of each table, we have included the average percentage changes.
@@
The average results differ significantly across the frequencies, averaging 28% for annual frequency, 36% for quarterly, and 48% for monthly. These discrepancies, although substantial, follow a pattern: the accuracy increases linearly by approximately 10% with each increase in frequency. Notably, the increase in frequency corresponds to a rise in the amount of data provided. For instance, the annual frequency typically has around 50 data points, quarterly frequency has about 100 observations, and monthly frequency has approximately 300. Similarly, the testing data consists of around 30, 80, and 240 observations for annual, quarterly, and monthly frequencies, respectively. This translates to an approximate forecasting horizon of 30 years for annual data, decreasing to 20 years for both quarterly and monthly data. The increase in the amount of data is significantly steeper than the improvement in accuracy. The impact of this steep increase on performance may be mitigated by the fact that actual forecasting periods are becoming shorter, as EDMS forecasts tend to perform better for longer horizons than for shorter ones, as stated in the method description. Therefore, it is reasonable to conclude that we observe a linear increase in overall performance.
@@
The results show that the EDMS significantly outperforms the EIMS for almost every data set. This is particularly evident in high-variability datasets such as CPI, where EDMS achieved significantly better results. CPI is influenced by monthly price changes, often driven by energy, food, and housing costs. These can be subject to shocks (e.g., oil price changes). According to CPI by categories, this dataset is highly granular, with 497 time series. Each category (e.g., food, transport, healthcare) behaves differently. EDMS’s ability to adapt to diverse patterns across categories gives it an edge, because EDMS’s ability to integrate over time helps to adapt those patterns. Similarly, for GDP (real and nominal), EDMS provides more accurate forecasts. The only exception was Employment (annual), where EIMS slightly outperformed EDMS, probably due to the smoother and more predictable nature of demographic trends which benefit from iterative method. Economic time series like interest rates and production often exhibit non-linear trends. For example, changes in interest rates may affect investment and employment with a delay, and these effects may not be proportional. EDMS is better in handling these behaviours because it integrates the smoothing process over time, allowing it to adapt to long-term trends and gradual changes more effectively, e.g. interest rates often follow long cycles influenced by monetary policy and global capital flows. Production and prices can be affected by supply chain innovations or external shocks (e.g. commodity shocks/unexpected tariffs). In contrast, unemployment rate is influenced by structural factors like labour market reforms or demographic changes. EDMS can accommodate these slow-moving shifts, while EIMS might miss them or misinterpret them as noise. Employment data at monthly frequency reflects both seasonal hiring patterns and long-term labour market trends. These patterns can be irregular due to holidays, school cycles, or policy changes.
@@
In Table 4, we present the results of computations for financial data, where each row represents a distinct index, although the models were trained on data from all indices. The table shows the average MAPE errors (4) for the EIMS and EDMS methodologies, as well as the values of $\Delta(\%)$. The total number of observations and the number of observations for testing are consistent across all indices, with 1041 observations and 720 testing observations. Additionally, we have included the average of percentage changes at the bottom of the table.
@@
Another reason for the disrupted pattern of performance improvement with the amount of data is the substantially different nature of financial data. Specifically, the data is daily, and the market is more liquid. Furthermore, there is considerable noise in the data and potential for speculation, and indices react strongly and promptly to various types of disruptions, such as tariffs (i.e., shocks).
Additionally, we compute the average percentage improvement across the entire sample of data sets, which is equal to 33.32%. This significant improvement in overall accuracy supports the assertion made in the description of EDMS. Specifically, the methodology not only leverages the strengths of Ensemble and DMS methods but is also highly influenced by their synergistic effect. The synergy can be inferred from the fact that applying DMS alone to EIMS would not be as beneficial. This is because DMS generally provides only slight improvements, and not in all cases, as demonstrated in previous works on DMS discussed in the literature review.
@@ 5.1 Results of comparative analysis
In light of the extensive comparative analysis conducted for the newly introduced Ensembled Direct Multi-Step (EDMS) methodology, we can confirm our theoretical claim. The average improvement across all datasets was 33.32% in favor of EDMS. This suggests that incorporating just a few retrainings into the iterative ensembling process yields significantly better accuracy, regardless of the frequency of considered economic indicators of financial indices. Indeed, the only plausible explanation for the observed improvement in forecast accuracy is the synergy between the Ensembling technique and DMS, resulting in an overall enhancement that surpasses the benefits of each method when applied individually.
The improvement in accuracy for macroeconomic indicators follows a clear trend based on data frequency, with annual data averaging around 28%, quarterly data at approximately 36%, and monthly data reaching about 48%. Notably, the increase in the amount of data with increase of frequencies is much steeper than the corresponding improvement in accuracy. The steepness of the data size impact may be mitigated by shorter forecasting periods for higher frequencies, as EDMS forecasts should demonstrate improved performance over longer horizons. Thus, a linear increase in overall performance is reasonable. From an economic perspective, considered time series represent key macroeconomic indicators that help policymakers assess and forecast the direction of the economy. Economic relationships are rarely linear or static. They evolve over time, influenced by policy decisions, expectations, and external shocks. EDMS’s architecture is designed to adapt and retain long-term dependencies, which mirrors how past crises, policy changes, or demand shocks continue to influence current economic behaviour. This memory aspect is important factor in economics, where historical context often shapes future expectations and actions. EDMS models are inherently designed at capturing these nonlinear and dynamic patterns.
A significant improvement of 20% is observed for financial data. However, this performance increase does not align with the pattern of improvement seen with increasing amount of data for macroeconomic data. This irregularity may arise from the use of a significantly smaller and more homogeneous sample of time series. Moreover, financial data is substantially different from economic data, notably it is more variable and less predictable in nature. Therefore, since model specification is more demanding, theoretically, EDMS should perform even better than EIMS. This fact strengthens the claim that size of time series sample was insufficient.
@@ 5.3 Potential applications
The proposed approach can be naturally applied to forecasting other macroeconomic variables as well as various types of financial data. Importantly, there is nothing in the methodology that restricts its use to macroeconomic or financial data. Therefore, it is worth exploring the applicability of the proposed approach in other domains, such as healthcare (e.g. disease outbreak prediction), environmental science (e.g. weather and climate modeling), transportation (e.g. traffic flow and delivery time estimation), energy (e.g. load and generation prediction, renewable energy generation prediction).
Based on our research, this methodology is particularly well-suited for improving medium- and long-term economic and financial forecasting. As such, it holds strong potential for use in institutions where accurate mid-long term forecasting is critical, including: central banks or other regulatory institutions, to support monetary policy decisions and promote economic stability, governments to inform fiscal policy, assess macroeconomic trends, and develop strategies for sustainable growth.. Additionally, commercial banks and other financial institutions can leverage EDMS to enhance risk assessment, improve asset management strategies, and optimize investment decisions. Investment firms and individual investors may also find value in using the methodology to predict market trends and identify emerging opportunities, thereby enabling more informed decision-making.

### v2

@@ 4 Comparisons on macroeconomic and financial data
We evaluate EDMS methodology against the EIMS approach to quantify the performance gains achieved through direct forecasting. In Subsection 4.1, we describe the data used, along with the process of data gathering and preparation. Subsection 4.2 explains the methodology employed in the comparative analysis, while the final Subsection 4.3 presents and discusses the results.
@@ 4.1 Data
Our comparative analysis utilizes carefully selected, common, and publicly available datasets. We chose public data for transparency and to ensure that our methodology is tested with data that is relatively similar to data potentially used in the practical applications. Moreover, to ensure the objectivity of our results, we employ a well-defined process of data selection and preparation, which can be broken down into the following stages:
@@
    Indicators selection - a set of common macroeconomic indicators is selected.
@@
    Source selection - for the selected indicators, we identify the public source with the highest quality of data, specifically focusing on methodologically consistent datasets across multiple countries for the same indicators. Furthermore, countries are selected based on two primary criteria: geographical diversity and the level of economic development, aiming to capture a representative cross-section of the global economy.
@@
    Trimming data set - for the forecasting implementation, it is essential that all time series have equal lengths. However, the time series in the dataset vary in length. We refrain from reducing all time series to the shortest one, as this would significantly decrease the size of the dataset and ultimately diminish the quality of model training. Instead, we remove only those time series with the smallest number of observations while simultaneously reducing the overall length of all time series in a manner that maximizes the number of data points in the entire dataset.
@@
    Exclude outliers - we eliminate time series that are highly irregular and cause model fit errors that are disproportionately higher than those of other time series in the specific dataset.
@@
For data description with list of specific countries included see Appendix A. Additionally, further data was selected from the Connect Platform, which is available through trusted data provider S&P Global Market Intelligence:
@@
For data description with list of specific countries included see Appendix A. In the CPI and Short Interest Rate datasets, we excluded Brazil due to a sharp increase in its values for some periods, which caused the forecast error to peak at approximately 50 times higher than the errors of the other countries. A similar situation occurred for Brazil and Peru in the Wholesale Producer Price Index, where the forecast error peak was around 10 times greater than that of other countries.
Additionally, a comprehensive set of Harmonized Index of Consumer Prices (HICP) indicators was sourced from Eurostat for EU countries. These indicators included numerous concepts, all rebased to 2015. For the list of concepts included see Appendix A. These indices were analysed over the period from December 2000 to December 2024, in accordance with the broader dataset constraints.
@@
  • FTSE 100 refers to a broader index of major companies listed on the London Stock Exchange, reflecting the performance of the UK market,
@@
  • Nikkei 225 monitors 225 large, publicly owned companies on the Tokyo Stock Exchange and serves as a primary indicator of Japan’s stock market performance,
@@
We used daily data from 2021.07.14 until 2025.07.14.
We divided each dataset into a training and testing sets. Since one of the main features of DMS is it’s ability to avoid the accumulation of errors in consecutive forecasts, a problem that often occurs in IMS, hence, the testing data set must be sufficiently large. The another reason for large testing set is that difference from IMS might be rather small in small forecasting horizons, as the first retraining is carried out after one year/month, and for that time differences in forecasts might still relatively small. After the second retraining, that is, in five years/months time, the difference should be more visible. On the other hand, there should be sufficient training data to properly train models. The balance was found to have around 10-15 years/months of trainig data and 20-25 years/months of testing data.
@@ 4.2 Comparison methodology
We measure the performance of the EDMS methodology by comparing it with the EIMS, as EIMS is a conventional forecasting methodology and serves as a natural benchmark. The comparison involves calculating the Mean Absolute Percentage Error (MAPE) for each time series in the datasets and take the average of that. This denoted as $\overline{mape}$. We selected MAPE because this metric is well-suited for comparing multiple time series and calculating an average metric. For instance, in the case of Mean Absolute Error (MAE), each time series may have values of different magnitudes, leading to considerable variation in MAE scores. Consequently, the contributions of some series to the overall average may distort the final value. In contrast, MAPE mitigates this issue since all errors are expressed on the same scale, that is, percentages.
@@
where $M$ represents the number of time series for a particular dataset, $H$ is the number of observations in the forecasting horizon, $y_{t+h}^{j}$ are the testing observations of the $j$-th time series, and $t+h$ is the time step at which forecasts are computed. Here $\widehat{y}_{t+h}^{j}$ can be either, EDMS or EIMS forecast.
Recall that the forecast using EIMS methodology at time step $t+h$ can be expressed by the following formula:
$\widehat{y}_{t+h}=\sum_{i}w_{t,i}\widetilde{y}_{t+h,i}(Y_{t}),$
where $w_{t,i}$ are the weights computed from the initial dataset, $h=1,2,\ldots,H$ are the steps in the horizon, and $\widetilde{y}_{t+h,i}$ are the forecasts at a time step one year ahead from time $t$ for the $i$-th model, which is trained on the data up to time $t$, denoted as $Y_{t}=(y_{1},y_{2},\ldots,y_{N})$.
@@ 4.3 Results
In Tables 1, 2 and 3 we present the results of our computations for economic data. We categorize the results into annual, quarterly, and monthly frequencies due to the substantial differences in the amount of training data used for each frequency, which significantly impacts the analysis results. In the tables, each row corresponds to a distinct dataset. In each row, in addition to the average MAPE errors (4) for the EIMS and EDMS methodologies and the values of $\Delta(\%)$, we also list the number of time series, the total number of observations, and the number of observations for testing. At the bottom of each table, we have included the average percentage changes.
@@
The average improvement differ significantly across frequencies, averaging 28% for annual frequency, 36% for quarterly, and 48% for monthly. These discrepancies, although substantial, follow a pattern: the accuracy increases linearly by approximately 10% with each increase in frequency. Notably, the increase in frequency corresponds to a rise in the amount of data provided. For instance, the annual frequency typically has around 50 data points, quarterly frequency has about 100 observations, and monthly frequency has approximately 300. Similarly, the testing data consists of around 30, 80, and 240 observations for annual, quarterly, and monthly frequencies, respectively. This translates to an approximate forecasting horizon of 30 years for annual data, decreasing to 20 years for both quarterly and monthly data. we can notice that the increase in the amount of data is significantly steeper than the improvement in accuracy. The impact of this steep increase on performance may be mitigated by the fact that actual forecasting periods are becoming shorter, as EDMS forecasts tend to perform better for longer horizons than for shorter ones. Therefore, it is reasonable to conclude that we observe a linear increase in overall performance.
@@
The results show that the EDMS significantly outperforms the EIMS for almost every data set. This is particularly evident in high-variability datasets such as CPI, where EDMS achieved significantly better results. CPI is influenced by monthly price changes, often driven by energy, food, and housing costs. These can be subject to shocks (e.g., oil price changes). Similarly, for GDP (real and nominal), EDMS provides more accurate forecasts. The only exception was Employment (annual), where EIMS slightly outperformed EDMS, probably due to the smoother and more predictable nature of demographic trends which benefit from iterative method. Economic time series like interest rates and production often exhibit non-linear trends. For example, changes in interest rates may affect investment and employment with a delay, and these effects may not be proportional. EDMS is better in handling these behaviours because it integrates the smoothing process over time, allowing it to adapt to long-term trends and gradual changes more effectively, e.g. interest rates often follow long cycles influenced by monetary policy and global capital flows. Production and prices can be affected by supply chain innovations or external shocks (e.g. commodity shocks/unexpected tariffs). In contrast, unemployment rate is influenced by structural factors like labour market reforms or demographic changes. EDMS can accommodate these slow-moving shifts, while EIMS might miss them or misinterpret them as noise. Employment data at monthly frequency reflects both seasonal hiring patterns and long-term labour market trends. These patterns can be irregular due to holidays, school cycles, or policy changes, so here also EDMS can be advantageous.
@@
In Table 4, we present the results of computations for financial data, where each row represents a distinct index, although the models were trained on data from all indices. Each table’s row shows the average MAPE errors (4) for the EIMS and EDMS methodologies, as well as the values of $\Delta(\%)$. The total number of observations and the number of observations for testing are consistent across all indices, with 1041 observations in total, 720 observations used for testing, and remaining for training. Additionally, we have included the average of percentage changes at the bottom of the table.
@@
Another reason for the disrupted pattern of performance improvement with the amount of data, is the substantially different nature of financial data. Specifically, the data is daily, and the market is more liquid. Furthermore, there is considerable noise in the data and potential for speculation, and indices react strongly and promptly to various types of disruptions, such as tariffs, i.e., shocks.
Additionally, we compute the average percentage improvement across the entire sample of data sets, which is equal to 33.32%. This significant improvement in overall accuracy supports the assertion made in the description of EDMS. Specifically, that the methodology not only leverages the strengths of Ensemble and DMS methods but is also highly influenced by their synergistic effect. The synergy can be inferred from the fact that applying DMS alone to EIMS would not be as beneficial. This is because DMS generally provides only slight improvements, and not in all cases, as demonstrated in previous works on DMS discussed in the literature review.
@@ 5.1 Results of comparative analysis
In light of the extensive comparative analysis conducted for the newly introduced Ensembled Direct Multi-Step (EDMS) methodology, we can confirm our theoretical claim. The average improvement across all datasets was 33.32% in favour of EDMS. This suggests that incorporating just a few retrainings into the iterative ensembling process yields significantly better accuracy, regardless of the frequency of considered economic indicators. Indeed, the only plausible explanation for the observed improvement in accuracy is the synergy between the Ensembling technique and DMS, resulting in an overall enhancement that surpasses the benefits of each method when applied individually.
The improvement for macroeconomic indicators follows a clear trend based on data frequency, with annual data averaging around 28%, quarterly data at approximately 36%, and monthly data reaching about 48%. Notably, the increase in the amount of data with increase of frequencies is much steeper than the corresponding improvement in accuracy. The steepness of the data size impact may be mitigated by shorter forecasting periods for higher frequencies, as EDMS forecasts should demonstrate improved performance over longer horizons. Thus, a linear increase in overall performance is reasonable.
From an economic perspective, considered time series represent key macroeconomic indicators that help policymakers assess and forecast the direction of the economy. Economic relationships are rarely linear or static. They evolve over time, influenced by policy decisions, expectations, and external shocks. EDMS’s architecture is designed to adapt and retain long-term dependencies, which mirrors how past crises, policy changes, or demand shocks continue to influence current economic behaviour. This memory aspect is important factor in economics, where historical context often shapes future expectations and actions. EDMS models are inherently designed at capturing these nonlinear and dynamic patterns.
A significant improvement of 20% is observed for financial data. However, this performance increase does not align with the pattern of improvement seen with increasing amount of data for macroeconomic data. This irregularity may arise from the use of a significantly smaller and more homogeneous sample of time series. Moreover, financial data is substantially different from economic data, notably it is more variable and less predictable in nature. Therefore, since model specification is more demanding, theoretically, EDMS should perform even better than EIMS. This fact strengthens the claim that size of time series sample might have been insufficient.
@@ 5.3 Potential applications
The proposed approach can be naturally applied to forecasting other macroeconomic variables as well as other types of financial data. Importantly, there is nothing in the methodology that restricts its use to macroeconomic or financial data. Therefore, it is worth exploring the applicability of the proposed approach in other domains, such as healthcare (e.g. disease outbreak prediction), environmental science (e.g. weather and climate modelling), transportation (e.g. traffic flow and delivery time estimation), energy (e.g. load and generation prediction, renewable energy generation prediction). Based on our research, this methodology is particularly well-suited for improving medium- and long-term economic and financial forecasting. As such, it holds strong potential for use in institutions where accurate mid-long term forecasting is critical, including: central banks or other regulatory institutions, to support monetary policy decisions and promote economic stability, governments to inform fiscal policy, assess macroeconomic trends, and develop strategies for sustainable growth. Additionally, commercial banks and other financial institutions can leverage EDMS to enhance risk assessment, improve asset management strategies, and optimize investment decisions. Investment firms and individual investors may also find value in using the methodology to predict market trends and identify emerging opportunities, thereby enabling more informed decision-making.